OpenStudy (anonymous):

find cos 0 if sin 0= -5/13 and tan 0>0

OpenStudy (ankit042):

cos ^2 = 1-sin^2 Now as tan >0 and sin <0 this is 3rd quad where cos is also <0

OpenStudy (anonymous):

okay so what do I next because I have the choices of -5/12 -13/5 -12/13 12/5

OpenStudy (jdoe0001):

\(\bf sin(\theta) = -\cfrac{5}{13} \implies \cfrac{\textit{opposite side}}{\textit{hypotenuse}} \implies \cfrac{b}{c}\)

OpenStudy (anonymous):

okay I understand that what do I do next

OpenStudy (jdoe0001):

well, you want cosine, so \(\bf cos(\theta) = \cfrac{\textit{adjacent side}}{\textit{hypotenuse}} \implies \cfrac{a}{c}\)

OpenStudy (ankit042):

put value of sin in the equation and solve you will get cos^2 =1-25/169 Now you can solve from here